OpenStudy (anonymous):

One thousand raffle tickets are sold for $1.00 each. One grand prize of $400 and two consolation prizes of $100 each will be awarded. Jeremy purchases one ticket. Find his expected value.

OpenStudy (anonymous):

expected value is what you win times probability you win it added up

OpenStudy (anonymous):

or you can simply imagine you bought all the tickets. you would spend 1000 and win one 400 prize and two 100 prizes for a grand total of 600. you spent 1000, you won 600 for a net loss of 400 averaged over the 1000 tickets this is -.4 or -40 cents per ticket

OpenStudy (anonymous):

ok that makes sense...

OpenStudy (anonymous):

in simple english he expects to lose 40 cents

OpenStudy (anonymous):

ok... i understand that.. thank you.. :)

OpenStudy (anonymous):

i hope your teacher or whomever doesn't want you to say the expected value is .60 that ignores the fact that you spent $1.00

OpenStudy (anonymous):

who knows. the prof dont explain anything... i am completely bombing this class.. :(

OpenStudy (anonymous):

we can write this using just probability if you like

OpenStudy (anonymous):

it amounts to the same thing exactly. the probability you lose your $1 is \[\frac{997}{1000}=.997\] the probability you win $399 is \[\frac{1}{1000}=.001\] (i wrote $399 because you do not get your dollar back too!) and the probability you win $99 is \[\frac{2}{1000}=.002\]

OpenStudy (anonymous):

expected value is \[-1\times .997+399\times .001+99\times .002=-.4\]

OpenStudy (anonymous):

basically its the same outcome.. just broke down in a different way

OpenStudy (anonymous):

yes first way you can almost do in your head. second way requires some calculation

OpenStudy (anonymous):

is it just a probability class?

OpenStudy (anonymous):

no its a college mathematics.. this week is about probability

OpenStudy (anonymous):

right one from column a and one from column b. chinese menu math nothing is worse

OpenStudy (anonymous):

it is a 6 week course and were in week 4 right now.. i think im catching on and then i end up bombing big time.. its getting frustrating.. big time.

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

thank you!